OpenStudy (megans20):

Amanda made $161 for 7 hours of work. At the same rate, how many hours would she have to work to make $229?

OpenStudy (megans20):

I got 13 hours but I'm not sure if my method is what my teacher is looking for.. I don't know what equation would solve this.

OpenStudy (kropot72):

The first step is to find how much she earns for each hour of work: \[\large Hourly\ rate=\frac{161}{7}=23\] So, she earns $23 per hour. Therefore is she worked for 10 hours she would earn 23 * 10 = $230. Therefore to earn exactly $229 she must work slightly less than 10 hours: \[\large \frac{229}{23}=9.96\ hours\]

OpenStudy (megans20):

Thank you so much! (:

OpenStudy (kropot72):

You're welcome :)